# Beyond Thumbs Up/Down: Untangling Challenges of Fine-Grained Feedback for Text-to-Image Generation

### v1

@@ Abstract.
text-to-image models, human feedback, efficient elicitation, reward modeling
journalyear: 2024copyright: noneconference: ; booktitle: ;11footnotetext: Work conducted while a Student Researcher at Google DeepMind22footnotetext: Work conducted while a Student Researcher at Google Research33footnotetext: Work conducted while at Google Research11footnotetext: Work conducted while at Google DeepMind
Figure 1. Example text-image pair where granular feedback matters.
Figure 2. Top: a typical coarse-grained feedback reward pipeline; bottom: proposed method for modeling fine-grained feedback.
1. Introduction
Human feedback serves as a critical element in adapting large-scale generative models, particularly within the Reinforcement Learning with Human Feedback (RLHF) paradigm (Casper et al., 2023; Christiano et al., 2017; Ouyang et al., 2022). However, conventional methods often rely on coarse-grained feedback, such as a single binary preference or Likert-scale rating, which may not adequately capture the nuances of quality and prompt-alignment in complex domains like text-to-image generation. A generated image may be highly visually appealing while deviating from the prompt, or conversely, align with the prompt and target visuals in some but not all desired ways, as illustrated in Figure 1. Recent research in text-to-text models (i.e. Large Language Models) suggests the potential of fine-grained feedback to address this challenge by enabling users to express their preferences with greater granularity, targeting specific features or interactions within the generated output (Wu et al., 2023a; Lee et al., 2023; Liang et al., 2023). This granular feedback mechanism promises more precise control over model adaptation and behavior, which could lead to text-to-image models that exhibit enhanced responsiveness to diverse user needs and preferences. For example, a user may want more photorealistic owls for a presentation but less for a fun t-shirt (Dutta et al., 2024).
@@ 2. Related Work
The rise of generative AI systems and power of RLHF – fueled by human feedback – has ignited further interest in gleaning insights from human data at scale to guide improved generation, e.g., to better align images to text prompts (Yarom et al., 2023; Hu et al., 2023; Kirstain et al., 2023; Xu et al., 2023) or to personalize image-generation models (von Rutte et al., 2023; Fan et al., 2023). Improved generative models offer immense potential to transform human productivity and creativity – provided they adequately meet the diverse needs and preferences of users. It is not yet clear what kind of feedback is best to elicit to improve their output – nor how best to incorporate such feedback when elicited. Works have taken steps to explore human feedback in text-to-image contexts  (Liang et al., 2023; Lee et al., 2023) as well as at evaluation-time  (Lee et al., 2024). Yet, the sheer computational scale of such generative models renders them challenging to systematically explore design choices around feedback elicitation and incorporation. Here, we take steps to address this gap by offering further empirical glimpses into the nuanced value of fine-grained feedback in generative AI applications.
@@
Several prior works have investigated the interplay between richer forms of human feedback and model performance. Notably, recent research in the text-to-text domain has explored the potential of fine-grained feedback (Wu et al., 2023a; Lee et al., 2023; Liang et al., 2023; Ouyang et al., 2022). In these studies, “fine-grained” refers to feedback that goes beyond simple binary judgments or single-score ratings, allowing users to target specific aspects of the output, such as factual accuracy, logical coherence, or stylistic elements. This granular feedback enables more precise control over model adaptation and behavior, leading to outputs that better align with diverse user preferences. In addition to granularity, richer feedback can also encompass representations of disagreement and uncertainty in human labels. For instance, researchers have explored the value of eliciting and learning with traces of human uncertainty in the form of soft labels (Peterson et al., 2019; Uma et al., 2020; Collins et al., 2022, 2023b; Sucholutsky et al., 2023).
@@
These studies have primarily focused on training classifiers for traditional machine learning tasks like image recognition, where human feedback is typically collected before training. However, practical applications often necessitate the ability to provide feedback on already trained models, potentially due to legislative requirements (e.g., Article 13 in the EU AI Act (European-Parliament, 2024)) or evolving user needs. Concept Bottleneck Models (CBMs), which map inputs to higher-level attributes before regressing a final target, offer a potential solution by allowing humans to provide feedback on a model’s intermediate outputs (Koh et al., 2020), which we explore in the remainder of our work for modeling fine-grained feedback.
3. Problem Setting
@@
Our goal is to learn a reward function $R_{\theta}:\mathcal{X}\rightarrow\mathcal{S}$ that takes in a set of features $x\in\mathcal{X}$ and produces a scalar score $s\in\mathcal{R}$ indicating the quality of $x$ [^2] The reward models we consider in this work produce pointwise, rather than pairwise, quality scores for each input.. Many popular reward learning-based frameworks exist that adapt pre-trained language (Ouyang et al., 2022) or diffusion-based models (Fan et al., 2024; Black et al., 2023; Dvijotham et al., 2023) to generate outputs with high reward scores.
@@
We consider the setting of eliciting fine-grained feedback where human annotators are asked to provide a set of $M$ scores $s_{i}=(s^{1}_{i},s^{2}_{i},...,s^{M}_{i})$ for each example $x_{i}$ with $i\in{1,\ldots,N}$. The individual scores are scalar values representing the degree to which the prompt-image pair satisfies some particular aspect of quality (e.g. photorealism). Our goal then is to learn a “good” (see Section 4) $R_{\theta}$ from such feedback. We refer to reward models learned from more than one feedback as “fine-grained reward models” compared to those trained on a single aggregate score, $\tilde{s}$ (“coarse-reward models”) and discuss different ways to concretely operationalize “good” in our experiments.
@@
While there seems to be intuitive value to collecting finer-grained feedback over more attributes, the elicitation of such feedback necessarily incurs some additional time (or other resource) cost, over the cost of collecting coarse-grained feedback. We consider the setting where each $j=1\ldots M$ dimension of feedback is associated with some elicitation cost $c_{j}>0$. For our computational experiments, we assume each form of feedback has equivalent procurement costs. We discuss deviations from this assumption in Section 6.2.1.
4. Reward Models from Fine-Grained Feedback
What kind of model structure empowers us to effectively learn from such rich feedback? We consider a two-stage structure which first predicts each fine-grained attributes and then aggregates the scores, similar to (Wu et al., 2023a). This structure parallels a Concept Bottleneck Model (Koh et al., 2020) (see above), wherein our concepts our fine-grained attributes (is the image malformed? is the image blurry? is the image aligned to the text for verbs?). Rather than learn a mapping directly from $R_{\theta}:x\rightarrow\tilde{s}$, we learn two sets of parameters: $f_{\phi}:x\rightarrow s^{1},...s^{M}$ and $g_{\psi}:s^{1},...s^{M}\rightarrow\tilde{s}$. The functions then compose to produce a single aggregate score for a given input $g_{\psi}(f_{\phi}(x))=\tilde{s}$ with the added benefit that we can inspect the fine-grained attributes predicted rendering our model more interpretable.
@@ 4. Reward Models from Fine-Grained Feedback
There are many functional forms that our two-stage modeling pipeline can take on. Here, we let $f$ be a multi-headed multi-layer perception (MLP). Following popular practice in the CBM literature, $g$ is a simple linear aggregator (Koh et al., 2020; Espinosa Zarlenga et al., 2022; Margeloiu et al., 2021; Collins et al., 2023a). We consider the sequential CBM setting (Koh et al., 2020), learning $\phi$ then $\psi$ separately.
@@
Additionally, in the text-to-image setting, we need some way of providing the multimodal stimulus (text and image) as input to our model. Following standard practice, we use learned embeddings of the image pixels and text tokens respectively. We assume that the feature extractors which produce the embeddings are fixed (see Appendix A.1). This approach has a benefit in common real-world scenarios where practitioners only have access to features in black-box fashion.
5. Experiment Outline
It is natural to expect that more informative supervision (finer-grained feedback) will be better for RLHF scenarios. In particular, we hypothesize that fine-grained feedback will be valuable when training reward models in regimes with few examples, following the literature around informativeness of label supervision (Sucholutsky et al., 2023). We posit based on previous results in text-to-text generation e.g. (Wu et al., 2023a) that reward models trained from fine-grained feedback will be able to better capture preference judgments than a model trained on coarse preference judgments alone.
To address our hypothesis, we design a series of computational experiments with feedback of varying levels of granularity and dataset sizes. We consider two classes of fine-grained feedback important for measuring the quality of a text-to-image generation following (Lee et al., 2024):
  1. (1)
@@ 5. Experiment Outline
Recall in our reward model that we employ a two-stage pipeline: first predicting fine-grained feedback, then predicting aggregate targets. As such, like in CBMs, we need two sets of labels (over fine-grained attributes and aggregate targets). In this work, the additional fine-grained feedback signals are inferred by querying the state-of-the-art PaLI model (Chen et al., 2022) (see Section 6) — we leave the expansion of collecting and incorporating granular feedback from human annotators for future work. In all experiments, the coarse- and fine-grained models are trained on the same final targets — real human preference judgments in Experiment 1 and a synthetic target permitting more controlled exploration in Experiment 2.
@@
6. Experiment 1: Predicting Human Preference Judgments from Fine-Grained Feedback
Figure 3. Comparing reward models trained on coarse feedback (i.e., direct human preference judgments; black) against CBM-based models learned from fine-grained feedback. Reward models are differentiated by whether they were trained on granular feedback only about image quality (blue), image-text alignment (red), or both (purple). Left: Each point represents a reward model trained on $N$ image-prompt examples (x axis); ROC-AUC of the binary reward against held-out human preference judgments is presented on the y axis. Higher is better. Right: The same reward models, where the x axis (presented on a log scale) depicts estimated annotation cost, if each attribute is assumed to be equally costly to procure.
@@ 6.1. Experimental Details
We use the approximately 5k images from  (Dutta et al., 2024). Images were generated from DALL-E  (Ramesh et al., 2021) and Stable Diffusion (Rombach et al., 2022) for over 1.3k text prompts from PartiPrompts  (Yu et al., 2022). Each image has been annotated by nine humans, where images are scored rated on a scale of 1-4 for how “satisfied” the viewer is with the image for the intended prompt, with a particular motivational context in mind (i.e., participants were asked to rate how good an image-prompt pair is for a particular motivational context: for use as a phone background, graphic t-shirt, or presentation slide-deck). Here, we average all context-conditioned scores to form a single preference (“goodness”) score per image-prompt pair. This score forms our “coarse”-grained preference of interest. We conduct a 50/25/25 train/val/test split at the level of the prompts (as there are four images per prompt). Additional details on data processing are included in the Appendix.
@@
The generative output of text-to-image models can be scored along both fidelity of the image to the prompt (text-image alignment (Yarom et al., 2023)) and image quality. We consider fine-grained attributes along each type: image quality and prompt alignment. For each image, we elicit granular quality attribute scores by querying PaLI (Chen et al., 2022) as discussed above. PaLI is a multimodal model which takes as input image and text and produces a text response; we can therefore use the model to simulate crowdsourced responses to text-image pairs by asking for a numerical score along some dimension. In particular, we train our model on eight image quality attributes (whether the image is distorted, photorealistic, bright, captivating, chaotic, visually compelling, disturbing, or funny). We select these attributes to span a range of valence and relevance. Additional details on granular queries are included in Appendix A.3.1.
To gather granular feedback for text-image alignment, we employ the $VQ^{2}$ framework from Yarom et al.. $VQ^{2}$ evaluates image-text alignment by generating question-answer pairs from the input text. For example, the text “A black apple and a green backpack” could yield the question-answer pair “What is the fruit in the picture? - Apple” or “What color is the backpack? - Green”. These pairs are then assessed against the image using a Visual Question Answering model, which produces a “Yes” probability indicating the validity of each pair. The average of these “Yes” probabilities across all question-answer pairs constitutes the final $VQ^{2}$ score, reflecting the overall image-text alignment. We group $VQ^{2}$ questions using a customized modular semantic parser into four attribute classes, whether the image matches the text in regards to: any actions/verbs mentioned in the prompt, any attributes/adjectives, any objects/nouns, or any relations mentioned in the prompt. See Appendix Section A.3.2 for further details on our feedback extraction.
@@
We compare a reward model trained directly on coarse feedback (the aggregate human scores from Dutta et al.) against a suite of CBMs trained on varying amounts ($d$) and types (e.g., image-based or prompt-based) fine-grained attributes. Modeling details are included in Appendix LABEL:model-details. Again, each attribute of fine-grained feedback acts like a “concept” predicted in the first stage of our two-stage model (Figure 2.
@@
  • Accuracy of the reward models, in terms of predicting the aggregated human “goodness” score for held-out examples from  (Dutta et al., 2024) (scored with ROC-AUC), and
@@ 6.2. Results
Interestingly, in Figure 3, we find that training reward models simply on coarse-grained feedback is not only more cost-efficient for a given budget (i.e., if we assume equal costs for collecting labels for each additional fine-grained attribute), but yield better fits than the putatively more “information-rich” fine-grained reward models. We discuss the impact of deviations from the equal-cost assumption in Section 6.2.1.
@@
6.2.1. Evaluation Proxy for Adapting a Generative Model with Reward Models
We take a step to address the first point by running a fresh human evaluation. Recall, one of our goals for learning reward models from fine-grained feedback is to better tailor adaptation of downstream generative models and improve the quality of their output rather than just matching aggregate human preference judgments – we made a case in our motivation that aggregate preference judgments can obfuscate important information. We simulate adapting a downstream generative model, and getting a sense of preference along the fine-grained attributes (focusing on the image quality attributes for simplicity). Due to computational costs, it is not always sensible to test out the quality of a gamut of reward models by adapting text-to-image models (see Section 8). Instead, we follow Lee et al. in running rejection sampling with our reward models as a proxy for adapting a generative model directly. That is, we draw samples from a generative model and use our reward models to score the outputs; we then run head-to-head preference judgments over the generations favored by the respective reward models. Here, we consider two reward models: the model trained only on coarse judgments (i.e., the aggregate human preference judgments from the train set) and our model trained on fine-grained simulated attribute annotations [^3] We consider our most “fine-grained”-trained model with $d=12$ attributes. Details on the sampled images and annotation procedures are included in the Appendix Section A.5.
We find in Table 1 (first row) that there is not a clear preference for the images that the reward model trained on coarse-grained feedback preferentially sample compared to the model trained on fine-grained feedback (i.e., annotators indicate that they prefer the text-image prompt rated more highly by the coarse-grained model than the fine-grained model for about 25% of the examples we survey, and vice versa for fine- over coarse-). This finding suggests that the difference between the coarse- and fine-grained feedback trained models are not as strong as our in-distribution prediction-based evaluation (from Sec. 6.2) make them appear as we move out-of-distribution to a new task: scoring generated images instead of judging (in-distribution) reward model accuracy. However, we observe high rates of uncertainty in the human judgements of which image is better along each attribute (annotators express that they are unsure which image they prefer for approximately 50% of the samples).
@@ 6.2.1. Evaluation Proxy for Adapting a Generative Model with Reward Models
Table 1. SxS preference scores, from rejection sampling. From these studies, we’ll get out a proportion of prompt-image pairs for which our raters prefer the example scored highly by the fine-grained model vs. the coarse-grained model (and vice versa), as well as the proportion where the annotators are unsure which is best. We compare pairwise general preferences (top row) as well as preference along particular granular attributes (rows below the line). Scores depict % of images where coarse- vs. fine-grained were preferred (or people were uncertain), where % depict the votes for each preferred image over the total number of votes.
Feedback Type Coarse Pref Fine-Grained Pref Unsure
@@
Table 2. Average annotator answer time (in seconds) for each annotation task.
@@
7. Experiment 2: Controlled Granular Image Quality Assessments
@@ 7.1. Experimental Setup
Figure 4. Comparing reward models trained on varying levels of granularity. As in Figure 3, each point represents a reward model trained on $N$ images. Models are scored according to the contrived decision tree on held-out examples. We compare a model trained directly on the single scalar decision tree scores (black) against a suite CBM-based fine-grained models trained on: 1) the same three attributes which make up the decision tree (red), 2) the same three attributes as the decision tree along with the remainder of the full set of image attributes under consideration (blue), and 3) only attributes not included in the decision tree (orange).
@@ 7.2. Results
Barring model architecture selection – we make a crucial assumption – that we know the true attributes. What if we do not have access to the attributes that form the decision tree? To begin to explore this question, we consider the setting where there is an attribute mismatch (yellow points in Fig 4). Here, we see a dramatic drop in performance. These data point to the importance of incorporating the right attributes if building a fine-grained reward model; here, we simulated and precisely controlled what attributes matter (by design of our decision tree). In practice, we may not know which attributes are “right” to elicit: the precise situation we found ourselves in for Experiment 1, potentially underlying our null result. One could envision selecting attributes that are most correlated with the target (see Figure 5), but this requires having access to target annotations. Determining what attributes to elicit is a key open challenge, as we discuss next.
Figure 5. Estimated similarity between PaLI scores for different attributes. We depict the proportion of images in the train set for which PaLI marks an image as having the same attribute (e.g., the cell blurry and malformed highlights that PaLI marks an image as blurry and malformed, or not blurry and not malformed for 71% of the examples). Darker red means higher level of similarity in scores; yellow represents lower similarity.
8. Open Challenges
@@
8.1. Which Attributes to Elicit?
@@
Our work points to a key challenge for practitioners interested in collecting fine-grained feedback: what attributes should you elicit? In Experiment 2, we demonstrate that reward model performance may suffer if the elicited attributes do not match those that form the target preferences. How can we know what attributes we should elicit? Such a question grows more challenging when we consider individual differences. Different attributes may matter to different people and depend on context (Gordon et al., 2022; Kirk et al., 2024b; Dutta et al., 2024). We do not want reward models to collapse to a monoculture (Kleinberg and Raghavan, 2021; Bommasani et al., 2022), but also ought to be mindful of the risks of personalization (Kirk et al., 2024a).

### v2

@@ Abstract
Figure 1: Example text-image pair where granular feedback matters.
Figure 2: Top: a typical coarse-grained feedback reward pipeline; bottom: proposed method for modeling fine-grained feedback.
Introduction
Human feedback serves as a critical element in adapting large-scale generative models, particularly within the Reinforcement Learning with Human Feedback (RLHF) paradigm (Casper et al. 2023; Christiano et al. 2017; Ouyang et al. 2022). However, conventional methods often rely on coarse-grained feedback, such as a single binary preference or Likert-scale rating, which may not adequately capture the nuances of quality and prompt-alignment in complex domains like text-to-image generation. A generated image may be highly visually appealing while deviating from the prompt, or conversely, align with the prompt and target visuals in some but not all desired ways, as illustrated in Figure 1. Recent research in text-to-text models (i.e. Large Language Models) suggests the potential of fine-grained feedback to address this challenge by enabling users to express their preferences with greater granularity, targeting specific features or interactions within the generated output (Wu et al. 2023b; Lee et al. 2023; Liang et al. 2023). This granular feedback mechanism promises more precise control over model adaptation and behavior, which could lead to text-to-image models that exhibit enhanced responsiveness to diverse user needs and preferences. For example, a user may want more photorealistic owls for a presentation but less for a fun t-shirt (Dutta et al. 2024).
@@ Introduction
The rise of generative AI systems and power of RLHF – fueled by human feedback – has ignited further interest in gleaning insights from human data at scale to guide improved generation, e.g., to better align images to text prompts (Yarom et al. 2023; Hu et al. 2023; Kirstain et al. 2023; Xu et al. 2023) or to personalize image-generation models (von Rutte et al. 2023; Fan et al. 2023). Improved generative models offer immense potential to transform human productivity and creativity – provided they adequately meet the diverse needs and preferences of users. It is not yet clear what kind of feedback is best to elicit to improve their output – nor how best to incorporate such feedback when elicited. Works have taken steps to explore human feedback in text-to-image contexts  (Liang et al. 2023; Lee et al. 2023) as well as at evaluation-time  (Lee et al. 2024). Yet, the sheer computational scale of such generative models renders them challenging to systematically explore design choices around feedback elicitation and incorporation. Here, we take steps to address this gap by offering further empirical glimpses into the nuanced value of fine-grained feedback in generative AI applications.
@@
Several prior works have investigated the interplay between richer forms of human feedback and model performance. Notably, recent research in the text-to-text domain has explored the potential of fine-grained feedback (Wu et al. 2023b; Lee et al. 2023; Liang et al. 2023; Ouyang et al. 2022). In these studies, “fine-grained” refers to feedback that goes beyond simple binary judgments or single-score ratings, allowing users to target specific aspects of the output, such as factual accuracy, logical coherence, or stylistic elements. This granular feedback enables more precise control over model adaptation and behavior, leading to outputs that better align with diverse user preferences. In addition to granularity, richer feedback can also encompass representations of disagreement and uncertainty in human labels. For instance, researchers have explored the value of eliciting and learning with traces of human uncertainty in the form of soft labels (Peterson et al. 2019; Uma et al. 2020; Collins, Bhatt, and Weller 2022; Collins et al. 2023b; Sucholutsky et al. 2023).
@@
These studies have primarily focused on training classifiers for traditional machine learning tasks like image recognition, where human feedback is typically collected before training. However, practical applications often necessitate the ability to provide feedback on already trained models, potentially due to legislative requirements (e.g., Article 13 in the EU AI Act (European-Parliament 2024)) or evolving user needs. Concept Bottleneck Models (CBMs), which map inputs to higher-level attributes before regressing a final target, offer a potential solution by allowing humans to provide feedback on a model’s intermediate outputs (Koh et al. 2020), which we explore in the remainder of our work for modeling fine-grained feedback.
Problem Setting
@@
Our goal is to learn a reward function $R_{\theta}:\mathcal{X}\rightarrow\mathcal{S}$ that takes in a set of features $x\in\mathcal{X}$ and produces a scalar score $s\in\mathcal{R}$ indicating the quality of $x$ [^1] The reward models we consider in this work produce pointwise, rather than pairwise, quality scores for each input.. Many popular reward learning-based frameworks exist that adapt pre-trained language (Ouyang et al. 2022) or diffusion-based models (Fan et al. 2024; Black et al. 2023; Dvijotham et al. 2023) to generate outputs with high reward scores.
@@
We consider the setting of eliciting fine-grained feedback where human annotators are asked to provide a set of $M$ scores $s_{i}=(s^{1}_{i},s^{2}_{i},...,s^{M}_{i})$ for each example $x_{i}$ with $i\in{1,\ldots,N}$. The individual scores are scalar values representing the degree to which the prompt-image pair satisfies some particular aspect of quality (e.g. photorealism). Our goal then is to learn a “good” (see Section Reward Models from Fine-Grained Feedback) $R_{\theta}$ from such feedback. We refer to reward models learned from more than one feedback as “fine-grained reward models” compared to those trained on a single aggregate score, $\tilde{s}$ (“coarse-reward models”) and discuss different ways to concretely operationalize “good” in our experiments.
@@
While there seems to be intuitive value to collecting finer-grained feedback over more attributes, the elicitation of such feedback necessarily incurs some additional time (or other resource) cost, over the cost of collecting coarse-grained feedback. We consider the setting where each $j=1\ldots M$ dimension of feedback is associated with some elicitation cost $c_{j}>0$. For our computational experiments, we assume each form of feedback has equivalent procurement costs. We discuss deviations from this assumption in Section Evaluation Proxy for Adapting a Generative Model with Reward Models.
Reward Models from Fine-Grained Feedback
What kind of model structure empowers us to effectively learn from such rich feedback? We consider a two-stage structure which first predicts each fine-grained attributes and then aggregates the scores, similar to (Wu et al. 2023b). This structure parallels a Concept Bottleneck Model (Koh et al. 2020) (see above), wherein our concepts our fine-grained attributes (is the image malformed? is the image blurry? is the image aligned to the text for verbs?). Rather than learn a mapping directly from $R_{\theta}:x\rightarrow\tilde{s}$, we learn two sets of parameters: $f_{\phi}:x\rightarrow s^{1},...s^{M}$ and $g_{\psi}:s^{1},...s^{M}\rightarrow\tilde{s}$. The functions then compose to produce a single aggregate score for a given input $g_{\psi}(f_{\phi}(x))=\tilde{s}$ with the added benefit that we can inspect the fine-grained attributes predicted rendering our model more interpretable.
@@
There are many functional forms that our two-stage modeling pipeline can take on. Here, we let $f$ be a multi-headed multi-layer perception (MLP). Following popular practice in the CBM literature, $g$ is a simple linear aggregator (Koh et al. 2020; Espinosa Zarlenga et al. 2022; Margeloiu et al. 2021; Collins et al. 2023a). We consider the sequential CBM setting (Koh et al. 2020), learning $\phi$ then $\psi$ separately.
@@
Additionally, in the text-to-image setting, we need some way of providing the multimodal stimulus (text and image) as input to our model. Following standard practice, we use learned embeddings of the image pixels and text tokens respectively. We assume that the feature extractors which produce the embeddings are fixed (see Appendix [^2] The full Appendix can be found in the arXiv version of this paper.). This approach has a benefit in common real-world scenarios where practitioners only have access to features in black-box fashion.
Experiment Outline
It is natural to expect that more informative supervision (finer-grained feedback) will be better for RLHF scenarios. In particular, we hypothesize that fine-grained feedback will be valuable when training reward models in regimes with few examples, following the literature around informativeness of label supervision (Sucholutsky et al. 2023). We posit based on previous results in text-to-text generation e.g. (Wu et al. 2023b) that reward models trained from fine-grained feedback will be able to better capture preference judgments than a model trained on coarse preference judgments alone.
To address our hypothesis, we design a series of computational experiments with feedback of varying levels of granularity and dataset sizes. We consider two classes of fine-grained feedback important for measuring the quality of a text-to-image generation following (Lee et al. 2024):
  1. 1.
@@
Recall in our reward model that we employ a two-stage pipeline: first predicting fine-grained feedback, then predicting aggregate targets. As such, like in CBMs, we need two sets of labels (over fine-grained attributes and aggregate targets). In this work, the additional fine-grained feedback signals are inferred by querying the state-of-the-art PaLI model (Chen et al. 2022) (see Section Experiment 1: Predicting Human Preference Judgments from Fine-Grained Feedback) — we leave the expansion of collecting and incorporating granular feedback from human annotators for future work. In all experiments, the coarse- and fine-grained models are trained on the same final targets — real human preference judgments in Experiment 1 and a synthetic target permitting more controlled exploration in Experiment 2.
@@
Experiment 1: Predicting Human Preference Judgments from Fine-Grained Feedback
Figure 3: Comparing reward models trained on coarse feedback (i.e., direct human preference judgments; black) against CBM-based models learned from fine-grained feedback. Reward models are differentiated by whether they were trained on granular feedback only about image quality (blue), image-text alignment (red), or both (purple). Left: Each point represents a reward model trained on $N$ image-prompt examples (x axis); ROC-AUC of the binary reward against held-out human preference judgments is presented on the y axis. Higher is better. Right: The same reward models, where the x axis (presented on a log scale) depicts estimated annotation cost, if each attribute is assumed to be equally costly to procure.
@@
We use the approximately 5k images from  (Dutta et al. 2024). Images were generated from DALL-E  (Ramesh et al. 2021) and Stable Diffusion (Rombach et al. 2022) for over 1.3k text prompts from PartiPrompts  (Yu et al. 2022). Each image has been annotated by nine humans, where images are scored rated on a scale of 1-4 for how “satisfied” the viewer is with the image for the intended prompt, with a particular motivational context in mind (i.e., participants were asked to rate how good an image-prompt pair is for a particular motivational context: for use as a phone background, graphic t-shirt, or presentation slide-deck). Here, we average all context-conditioned scores to form a single preference (“goodness”) score per image-prompt pair. This score forms our “coarse”-grained preference of interest. We conduct a 50/25/25 train/val/test split at the level of the prompts (as there are four images per prompt). Additional details on data processing are included in the Appendix.
@@
The generative output of text-to-image models can be scored along both fidelity of the image to the prompt (text-image alignment (Yarom et al. 2023)) and image quality. We consider fine-grained attributes along each type: image quality and prompt alignment. For each image, we elicit granular quality attribute scores by querying PaLI (Chen et al. 2022) as discussed above. PaLI is a multimodal model which takes as input image and text and produces a text response; we can therefore use the model to simulate crowdsourced responses to text-image pairs by asking for a numerical score along some dimension. In particular, we train our model on eight image quality attributes (whether the image is distorted, photorealistic, bright, captivating, chaotic, visually compelling, disturbing, or funny). We select these attributes to span a range of valence and relevance. Additional details on granular queries are included in the Appendix.
To gather granular feedback for text-image alignment, we employ the $VQ^{2}$ framework from Yarom et al.. $VQ^{2}$ evaluates image-text alignment by generating question-answer pairs from the input text. For example, the text “A black apple and a green backpack” could yield the question-answer pair “What is the fruit in the picture? - Apple” or “What color is the backpack? - Green”. These pairs are then assessed against the image using a Visual Question Answering model, which produces a “Yes” probability indicating the validity of each pair. The average of these “Yes” probabilities across all question-answer pairs constitutes the final $VQ^{2}$ score, reflecting the overall image-text alignment. We group $VQ^{2}$ questions using a customized modular semantic parser into four attribute classes, whether the image matches the text in regards to: any actions/verbs mentioned in the prompt, any attributes/adjectives, any objects/nouns, or any relations mentioned in the prompt. See Appendix for further details on our feedback extraction.
@@
We compare a reward model trained directly on coarse feedback (the aggregate human scores from Dutta et al.) against a suite of CBMs trained on varying amounts ($d$) and types (e.g., image-based or prompt-based) fine-grained attributes. Modeling details are included in the Appendix. Again, each attribute of fine-grained feedback acts like a “concept” predicted in the first stage of our two-stage model (Figure 2.
@@
  • Accuracy of the reward models, in terms of predicting the aggregated human “goodness” score for held-out examples from  (Dutta et al. 2024) (scored with ROC-AUC), and
@@ Results
Interestingly, in Figure 3, we find that training reward models simply on coarse-grained feedback is not only more cost-efficient for a given budget (i.e., if we assume equal costs for collecting labels for each additional fine-grained attribute), but yield better fits than the putatively more “information-rich” fine-grained reward models. We discuss the impact of deviations from the equal-cost assumption in Section Evaluation Proxy for Adapting a Generative Model with Reward Models.
@@
Evaluation Proxy for Adapting a Generative Model with Reward Models
We take a step to address the first point by running a fresh human evaluation. Recall, one of our goals for learning reward models from fine-grained feedback is to better tailor adaptation of downstream generative models and improve the quality of their output rather than just matching aggregate human preference judgments – we made a case in our motivation that aggregate preference judgments can obfuscate important information. We simulate adapting a downstream generative model, and getting a sense of preference along the fine-grained attributes (focusing on the image quality attributes for simplicity). Due to computational costs, it is not always sensible to test out the quality of a gamut of reward models by adapting text-to-image models (see Section Open Challenges). Instead, we follow Lee et al. in running rejection sampling with our reward models as a proxy for adapting a generative model directly. That is, we draw samples from a generative model and use our reward models to score the outputs; we then run head-to-head preference judgments over the generations favored by the respective reward models. Here, we consider two reward models: the model trained only on coarse judgments (i.e., the aggregate human preference judgments from the train set) and our model trained on fine-grained simulated attribute annotations [^3] We consider our most “fine-grained”-trained model with $d=12$ attributes. Details on the sampled images and annotation procedures are included in the Appendix.
We find in Table 1 (first row) that there is not a clear preference for the images that the reward model trained on coarse-grained feedback preferentially sample compared to the model trained on fine-grained feedback (i.e., annotators indicate that they prefer the text-image prompt rated more highly by the coarse-grained model than the fine-grained model for about 25% of the examples we survey, and vice versa for fine- over coarse-). This finding suggests that the difference between the coarse- and fine-grained feedback trained models are not as strong as our in-distribution prediction-based evaluation (from Sec. Results) make them appear as we move out-of-distribution to a new task: scoring generated images instead of judging (in-distribution) reward model accuracy. However, we observe high rates of uncertainty in the human judgements of which image is better along each attribute (annotators express that they are unsure which image they prefer for approximately 50% of the samples).
@@
Feedback Type Coarse Fine-Grained Unsure
@@
Table 1: We compare pairwise general preferences (top row) as well as preference along particular granular attributes (rows below the line). Scores depict % of images where coarse- vs. fine-grained were preferred (or people were uncertain), where % depict the votes for each preferred image over the total number of votes.
@@
Table 2: Average annotator answer time (in seconds) for each annotation task.
Experiment 2: Controlled Granular Image Quality Assessments
@@
Figure 4: Comparing reward models trained on varying levels of granularity. As in Figure 3, each point represents a reward model trained on $N$ images. Models are scored according to the contrived decision tree on held-out examples. We compare a model trained directly on the single scalar decision tree scores (black) against a suite CBM-based fine-grained models trained on: 1) the same three attributes which make up the decision tree (red), 2) the same three attributes as the decision tree along with the remainder of the full set of image attributes under consideration (blue), and 3) only attributes not included in the decision tree (orange).
@@
Barring model architecture selection – we make a crucial assumption – that we know the true attributes. What if we do not have access to the attributes that form the decision tree? To begin to explore this question, we consider the setting where there is an attribute mismatch (yellow points in Fig 4). Here, we see a dramatic drop in performance. These data point to the importance of incorporating the right attributes if building a fine-grained reward model; here, we simulated and precisely controlled what attributes matter (by design of our decision tree). In practice, we may not know which attributes are “right” to elicit: the precise situation we found ourselves in for Experiment 1, potentially underlying our null result. One could envision selecting attributes that are most correlated with the target (see Appendix), but this requires having access to target annotations. Determining what attributes to elicit is a key open challenge, as we discuss next.
Open Challenges
@@
Which Attributes to Elicit?
@@
Our work points to a key challenge for practitioners interested in collecting fine-grained feedback: what attributes should you elicit? In Experiment 2, we demonstrate that reward model performance may suffer if the elicited attributes do not match those that form the target preferences. How can we know what attributes we should elicit? Such a question grows more challenging when we consider individual differences. Different attributes may matter to different people and depend on context (Gordon et al. 2022; Kirk et al. 2024b; Dutta et al. 2024). We do not want reward models to collapse to a monoculture (Kleinberg and Raghavan 2021; Bommasani et al. 2022), but also ought to be mindful of the risks of personalization (Kirk et al. 2024a).